OpenStudy (anonymous):

Suppose you roll one die. What is the probability of rolling a multiple of 3?

OpenStudy (anonymous):

considering it is a fair dice, n(a) it'd be 3 and 6 only therefore n(s) = 6 P(a) = 2/6 = 1/3 therefore its 1/3